OpenStudy (anonymous):

Solve for x... 8x+1=8(x+9)

OpenStudy (kl0723):

so from the original equation you have, the left side seems to be ok, but what about the right side? can you distribute the 8 into (x+9)?

OpenStudy (anonymous):

8x+72

OpenStudy (kl0723):

great, so now you have the following equation: 8x + 1 = 8x + 72

OpenStudy (kl0723):

so now when you gather like terms you have 8x - 8x = 72 - 1

OpenStudy (kl0723):

oops, sorry my connection sucks...

OpenStudy (kl0723):

so you have now that the x's cancel out and that you have 72 = 1, is that true?

OpenStudy (anonymous):

no, so there's no solution, thanks :)

OpenStudy (kl0723):

ok :)